# Small time asymptotics of spectral heat content of isotropic processes

### v1

@@ 5. Proof of the lower bound
$\liminf_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}\geqslant\mathrm{Per}(\Omega).$
@@
When $\Omega$ has $C^{1,1}$-boundary, the above proposition can be proved using a half-space argument similar to the proof of [9, Lemma 3.1]. However, the argument in the aforementioned paper fails when $\partial\Omega$ is not regular.
@@
$\liminf_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}\geqslant\liminf_{t\to 0}\frac{\mathcal{R}^{X}_{f_{\varepsilon}}(t)}{\mu(t)}=\int_{\mathbb{R}^{d}}|\nabla f_{\varepsilon}(x)|dx.$
@@
$\liminf_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}\geqslant\liminf_{\varepsilon\to 0}\int_{\mathbb{R}^{d}}|\nabla f_{\varepsilon}(x)|\mu(dx)\geqslant\|D\mathbbm{1}_{\Omega}\|=\mathrm{Per}(\Omega).$
@@
$\mathrm{Per}(\Omega)\leqslant\liminf_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}\leqslant\limsup_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}\leqslant\frac{\sup_{\partial\Omega}|\nabla\phi|}{\inf_{\partial\Omega}|\nabla\phi|}\mathrm{Per}(\Omega).$
@@
$\lim_{t\to 0}\frac{1}{\mu(t)}\int_{\Omega_{\varepsilon}}\mathbb{P}_{x}\left(\inf_{0\leqslant s\leqslant t}\phi(X_{s})\leqslant 0\right)dx=0$
@@
$\ \ \ \ \lim_{t\to 0}\frac{1}{\mu(t)}\int_{\Omega_{\varepsilon}}\mathbb{P}_{x}\left(\inf_{0\leqslant s\leqslant t}\phi(X_{s})\leqslant 0\right)dx$
$\leqslant\lim_{t\to 0}\frac{1}{\mu(t)}\left(\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}|X_{s}|>\frac{\varepsilon}{2M}\right)+\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}|X_{s}|>\left(\frac{\varepsilon}{L^{\prime}}\right)^{1/1+\kappa}\right)\right)\mathrm{Vol}(\Omega)$
@@
$\limsup_{t\to 0}\frac{1}{\mu(t)}\int_{\Omega^{\prime}_{\varepsilon}}\mathbb{P}_{x}\left(\inf_{0\leqslant s\leqslant t}\phi(X_{s})\leqslant 0\right)dx$
@@
$=\int_{0}^{\varepsilon}\int_{\partial\Omega_{v}}\mathbb{P}_{0}\left(|\nabla\phi(x)|\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}>\eta v\right)\frac{\mathcal{H}^{d-1}(dx)}{|\nabla\phi(x)|}dv.$
@@
$\leqslant\sup_{0\leqslant v\leqslant\varepsilon}\mathcal{H}^{d-1}(\partial\Omega_{v})\frac{1}{m_{\varepsilon}}\int_{0}^{\varepsilon}\mathbb{P}_{0}\left(M_{\varepsilon}\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}>\eta v\right)dv$
$=\sup_{0\leqslant v\leqslant\varepsilon}\mathcal{H}^{d-1}(\partial\Omega_{v})\frac{M_{\varepsilon}}{\eta m_{\varepsilon}}\mathbb{E}_{0}\left(\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\wedge\frac{\eta\varepsilon}{M_{\varepsilon}}\right)$
@@
(6.5) $\limsup_{t\to 0}\frac{1}{\mu(t)}\int_{\Omega^{\prime}_{\varepsilon}}\mathbb{P}_{x}\left(\sup_{0\leqslant s\leqslant t}\langle X_{s}-x,\nabla\phi(x)\rangle>\eta\phi(x)\right)dx$
@@
$\leqslant((1-\eta)\varepsilon)^{\frac{1}{\kappa+1}}\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}|X_{s}|\wedge(1-\eta)^{\frac{1}{1+\kappa}}\varepsilon^{\frac{1}{1+\kappa}}\right]\leqslant 2\mu(t)((1-\eta)\varepsilon)^{\frac{\kappa}{\kappa+1}}$
@@
(6.7) $\limsup_{t\to 0}\frac{1}{\mu(t)}\int_{\Omega^{\prime}_{\varepsilon}}\mathbb{P}_{x}\left(\sup_{0\leqslant s\leqslant t}|X_{s}-x|^{1+\kappa}>(1-\eta)\phi(x)\right)dx$
@@
$\limsup_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}\leqslant\frac{\sup_{\partial\Omega}|\nabla\phi|}{\inf_{\partial\Omega}|\nabla\phi|}\mathrm{Per}(\Omega).$
@@
(6.8) $\ \ \ \ \limsup_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}=\int_{\Omega}\mathbb{P}_{x}(\tau^{X}_{\Omega}\leqslant t)dx$
$\leqslant\limsup_{t\to 0}\frac{1}{\mu(t)}\int_{\Omega_{\varepsilon}}\mathbb{P}_{x}\left(\inf_{0\leqslant s\leqslant t}\phi(X_{s})\leqslant 0\right)dx$
$+\limsup_{t\to 0}\frac{1}{\mu(t)}\int_{\Omega^{\prime}_{\varepsilon}}\mathbb{P}_{x}\left(\inf_{0\leqslant s\leqslant t}\phi(X_{s})\leqslant 0\right)dx.$
@@
$\ \ \ \ \limsup_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}$
@@
$\limsup_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}\leqslant\frac{\sup_{\partial\Omega}|\nabla\phi|}{\inf_{\partial\Omega}|\nabla\phi|}\mathcal{H}^{d-1}(\partial\Omega)=\frac{\sup_{\partial\Omega}|\nabla\phi|}{\inf_{\partial\Omega}|\nabla\phi|}\mathrm{Per}(\Omega).$
@@
The final step of the proof of Theorem 1.1 is to construct a function $\phi$ satisfying conditions of Theorem 6.1 for domains with $C^{1,1}$-boundary.
@@
Let $\Omega\subset\mathbb{R}^{d}$ be a bounded open set with $C^{1,1}$-boundary. Then there exists a bounded $C^{1,1}$ function $\phi:\mathbb{R}^{d}\to\mathbb{R}$ such that
@@
Since $\Omega$ has $C^{1,1}$-boundary, there exists $r>0$ such that $\delta$ is $C^{1,1}$ in $\Omega^{\prime}_{r}:=\{x\in\mathbb{R}^{d}:|\delta(x)|<r\}$ and $|\nabla\delta|=1$ in $\Omega^{\prime}_{r}$. Define a $C^{1,1}$ function $\varphi:\mathbb{R}\to\mathbb{R}$ such that $\varphi(x)=x$ when $|x|<r/2$ and equals $r$ when $|x|>r$ and consider $\phi=\varphi\circ\delta$. Clearly, $\phi$ is $C^{1,1}$ and $\phi(x)>0$ if and only if $x\in\Omega$. Also, for any $x\in\Omega^{\prime}_{r/2}$, $|\nabla\phi(x)|=|\varphi^{\prime}(\delta(x))\nabla\delta(x)|=1$. This completes the proof of the lemma. ∎
@@
The $C^{1,1}$-regularity assumption on $\partial\Omega$ cannot be weakened. From [8, Corollary 6] it is known that the differentiability of $\delta$ in a neighborhood of $\partial\Omega$ implies that $\partial\Omega$ is $C^{1,1}$.
@@
Proof of Corollary 1.3.
For any $\varepsilon>0$, using Markov inequality we obtain
$\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}>\varepsilon\right)\leqslant\frac{\mathbb{E}_{0}\left[\left(\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\right)^{\alpha}\right]}{\varepsilon^{\alpha}}.$
Also, from the given conditions and noting that $\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\geqslant 0$ if $X_{0}=0$, we have
$0\leqslant 1-\frac{\mu(t)}{m(t)}\leqslant\frac{\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\mathbbm{1}\left\{\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}>1\right\}\right]}{m(t)}\leqslant\frac{\mathbb{E}_{0}\left[\left(\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\right)^{\alpha}\right]}{m(t)},$
and the right hand side goes to 0 as $t\to 0$. Hence, $\mu(t)/m(t)\to 1$ as $t\to 0$ and Assumption 1.1 is satisfied. The proof is concluded by applying Theorem 1.1. ∎

### v3

@@ 5. Proof of the lower bound
$\liminf_{t\downarrow 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}\geqslant\mathrm{Per}(\Omega).$
@@
When $\Omega$ has $C^{1,1}$ boundary, the above proposition can be proved using a half-space argument similar to the proof of [18, Lemma 3.1]. However, the argument in the aforementioned paper requires the $R$-smooth boundary condition, which does not hold when $\partial\Omega$ is not $C^{1,1}$.
@@
$\liminf_{t\downarrow 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}\geqslant\liminf_{t\downarrow 0}\frac{\mathcal{R}^{X}_{f_{\varepsilon}}(t)}{\mu(t)}=\int_{\mathbb{R}^{d}}|\nabla f_{\varepsilon}(x)|dx.$
@@
$\liminf_{t\downarrow 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}\geqslant\liminf_{\varepsilon\to 0}\int_{\mathbb{R}^{d}}|\nabla f_{\varepsilon}(x)|dx\geqslant\|D\mathbbm{1}_{\Omega}\|=\mathrm{Per}(\Omega).$
@@
$\mathrm{Per}(\Omega)\leqslant\liminf_{t\downarrow 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}\leqslant\limsup_{t\downarrow 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}\leqslant\frac{\sup_{\partial\Omega}|\nabla\phi|}{\inf_{\partial\Omega}|\nabla\phi|}\mathrm{Per}(\Omega).$
@@
$\lim_{t\downarrow 0}\frac{1}{\mu(t)}\int_{\Omega_{\varepsilon}}\mathbb{P}_{x}\left(\inf_{0\leqslant s\leqslant t}\phi(X_{s})\leqslant 0\right)dx=0$
@@
$\ \ \ \ \lim_{t\downarrow 0}\frac{1}{\mu(t)}\int_{\Omega_{\varepsilon}}\mathbb{P}_{x}\left(\inf_{0\leqslant s\leqslant t}\phi(X_{s})\leqslant 0\right)dx$
$\leqslant\lim_{t\downarrow 0}\frac{1}{\mu(t)}\left(\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}|X_{s}|>\frac{\varepsilon}{2M}\right)+\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}|X_{s}|>\left(\frac{\varepsilon}{L^{\prime}}\right)^{1/1+\kappa}\right)\right)\mathrm{Vol}(\Omega)$
@@
$\limsup_{t\downarrow 0}\frac{1}{\mu(t)}\int_{\Omega^{\prime}_{\varepsilon}}\mathbb{P}_{x}\left(\inf_{0\leqslant s\leqslant t}\phi(X_{s})\leqslant 0\right)dx$
@@
$=\int_{0}^{\varepsilon}\int_{\partial\Omega_{v}}\mathbb{P}_{0}\left(|\nabla\phi(x)|\overline{X}^{(1)}_{t}>\eta v\right)\frac{\mathcal{H}^{d-1}(dx)}{|\nabla\phi(x)|}dv.$
@@
$\leqslant\sup_{0\leqslant v\leqslant\varepsilon}\mathcal{H}^{d-1}(\partial\Omega_{v})\frac{1}{m_{\varepsilon}}\int_{0}^{\varepsilon}\mathbb{P}_{0}\left(M_{\varepsilon}\overline{X}^{(1)}_{t}>\eta v\right)dv$
$=\sup_{0\leqslant v\leqslant\varepsilon}\mathcal{H}^{d-1}(\partial\Omega_{v})\frac{M_{\varepsilon}}{\eta m_{\varepsilon}}\mathbb{E}_{0}\left(\overline{X}^{(1)}_{t}\wedge\frac{\eta\varepsilon}{M_{\varepsilon}}\right)$
@@
(6.5) $\limsup_{t\downarrow 0}\frac{1}{\mu(t)}\int_{\Omega^{\prime}_{\varepsilon}}\mathbb{P}_{x}\left(\sup_{0\leqslant s\leqslant t}\langle X_{s}-x,\nabla\phi(x)\rangle>\eta\phi(x)\right)dx$
@@
$\leqslant((1-\eta)\varepsilon)^{\frac{\kappa}{\kappa+1}}\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}|X_{s}|\wedge(1-\eta)^{\frac{1}{1+\kappa}}\varepsilon^{\frac{1}{1+\kappa}}\right]\leqslant 2\mu(t)((1-\eta)\varepsilon)^{\frac{\kappa}{\kappa+1}}$
@@
(6.7) $\limsup_{t\downarrow 0}\frac{1}{\mu(t)}\int_{\Omega^{\prime}_{\varepsilon}}\mathbb{P}_{x}\left(\sup_{0\leqslant s\leqslant t}|X_{s}-x|^{1+\kappa}>(1-\eta)\phi(x)\right)dx$
@@
$\limsup_{t\downarrow 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}\leqslant\frac{\sup_{\partial\Omega}|\nabla\phi|}{\inf_{\partial\Omega}|\nabla\phi|}\mathrm{Per}(\Omega).$
@@
(6.8) $\ \ \ \ \limsup_{t\downarrow 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}=\int_{\Omega}\mathbb{P}_{x}(\tau^{X}_{\Omega}\leqslant t)dx$
$\leqslant\limsup_{t\downarrow 0}\frac{1}{\mu(t)}\int_{\Omega_{\varepsilon}}\mathbb{P}_{x}\left(\inf_{0\leqslant s\leqslant t}\phi(X_{s})\leqslant 0\right)dx$
$+\limsup_{t\downarrow 0}\frac{1}{\mu(t)}\int_{\Omega^{\prime}_{\varepsilon}}\mathbb{P}_{x}\left(\inf_{0\leqslant s\leqslant t}\phi(X_{s})\leqslant 0\right)dx.$
@@
$\ \ \ \ \limsup_{t\downarrow 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}$
@@
$\limsup_{t\downarrow 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}\leqslant\frac{\sup_{\partial\Omega}|\nabla\phi|}{\inf_{\partial\Omega}|\nabla\phi|}\mathcal{H}^{d-1}(\partial\Omega)=\frac{\sup_{\partial\Omega}|\nabla\phi|}{\inf_{\partial\Omega}|\nabla\phi|}\mathrm{Per}(\Omega).$
@@
The final step of the proof of Theorem 1.1 is to construct a function $\phi$ satisfying conditions of Theorem 6.1 for domains with $C^{1,1}$ boundary.
@@
Let $\Omega\subset\mathbb{R}^{d}$ be a bounded open set with $C^{1,1}$ boundary. Then there exists a bounded $C^{1,1}$ function $\phi:\mathbb{R}^{d}\to\mathbb{R}$ such that
@@
Since $\Omega$ has $C^{1,1}$ boundary, there exists $r_{0}>0$ such that $\delta$ is $C^{1,1}$ in $\Omega^{\prime}_{r_{0}}:=\{x\in\mathbb{R}^{d}:|\delta(x)|<r_{0}\}$ and $|\nabla\delta|=1$ in $\Omega^{\prime}_{r_{0}}$. Fix $0<r<r_{0}$ and define a $C^{1,1}$ function $\varphi:\mathbb{R}\to\mathbb{R}$
$\varphi(x)=\begin{cases}x&\mbox{if $|x|<r/2$}\\ r&\mbox{if $x\geqslant r$}\\ -r&\mbox{if $x\leqslant-r$}.\end{cases}$
Consider the function $\phi=\varphi\circ\delta$. As $\delta$ is $C^{1,1}$ in $\Omega^{\prime}_{r_{0}}$ and $\phi$ is constant on $\{x:\delta(x)\geqslant r\}$, $\phi$ is a bounded $C^{1,1}$ function on $\mathbb{R}^{d}$. Also, $\phi(x)>0$ if and only if $x\in\Omega$. For any $x\in\Omega^{\prime}_{r/2}$, $|\nabla\phi(x)|=|\varphi^{\prime}(\delta(x))\nabla\delta(x)|=1$. This completes the proof of the lemma. ∎
@@
The $C^{1,1}$ regularity assumption on $\partial\Omega$ cannot be weakened. From [17, Corollary 6] it is known that the differentiability of $\delta$ in a neighborhood of $\partial\Omega$ implies that $\partial\Omega$ is $C^{1,1}$.
@@
Proof of Proposition 1.3.
Since $\mathbb{P}_{0}(\overline{X}^{(1)}_{t}\geqslant 0)=1$, for any $\varepsilon>0$, using Markov inequality we obtain
$\mathbb{P}_{0}\left(\overline{X}^{(1)}_{t}>\varepsilon\right)\leqslant\frac{\mathbb{E}_{0}\left[\left(\overline{X}^{(1)}_{t}\right)^{p}\right]}{\varepsilon^{p}}.$
Also, from the given conditions we have
$0\leqslant 1-\frac{\mu(t)}{m(t)}\leqslant\frac{\mathbb{E}_{0}\left[\overline{X}^{(1)}_{t}\mathbbm{1}\left\{\overline{X}^{(1)}_{t}>1\right\}\right]}{m(t)}\leqslant\frac{\mathbb{E}_{0}\left[\left(\overline{X}^{(1)}_{t}\right)^{p}\right]}{m(t)},$
and the right hand side goes to 0 as $t\downarrow 0$, where $m(t)=\mathbb{E}_{0}[\overline{X}^{(1)}_{t}$]. Hence, $\mu(t)/m(t)\to 1$ as $t\downarrow 0$ and Assumption 1.1 is satisfied. The proof is concluded by applying Theorem 1.1. ∎